OpenStudy (anonymous):

FAN A set of equations is given below: Equation A: e = 4f + 1 Equation B: e = 3f + 5 Which of the following is a step that can be used to find the solution to the set of equations?

OpenStudy (anonymous):

@DaBest21

OpenStudy (anonymous):

Since e=e and e=4f+1 and e=3f+5 could u say 4f+1=3f+5 @hayespurdy

OpenStudy (anonymous):

what are the answer choices @hayespurdy

OpenStudy (anonymous):

4f + 1= 3f 4f = 3f + 5 4f + 5 = 3f + 1 4f + 1 = 3f + 5

OpenStudy (anonymous):

Meg is 6 years older than Victor. Meg's age is 2 years less than five times Victor's age. The equations below model the relationship between Meg's age (m) and Victor's age (v): m = v + 6 m = 5v − 2 Which is a possible correct method to find Meg's and Victor's ages? Solve m + 6 = 5m − 2 to find the value of m. Write the points where the graphs of the equations intersect the x axis. Solve v + 6 = 5v − 2 to find the value of v. Write the points where the graphs of the equations intersect the y axis. what about dis @DaBest21